OpenStudy (anonymous):

Which of the following is equivalent to 5 4/9 divided by 7 ? 7/9 1 2/7 5/7 5 4/63

OpenStudy (anonymous):

first convert it to an improper fraction: 5 4/9 = 49/9 then divide it by 7, which is just multiplying by the reciprocal 49/9 x 1/7 = 7/9